OpenStudy (anonymous):

how the expanded form and simplification of x^6 ---- x^2 . Explain in your own words how you can simplify x^6 without having to write the expanded form. --- x^2 Part 2: Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (anonymous):

Explain in your own words how you can simplify x^6 without having to write the expanded form. x^6 --- x^2

jimthompson5910 (jim_thompson5910):

Shortcut: \[\Large \frac{x^6}{x^2}\] \[\Large x^{6-2}\] \[\Large x^{4}\] So \[\Large \frac{x^6}{x^2}=x^{4}\]

jimthompson5910 (jim_thompson5910):

Basically, if you divide variable terms, you subtract corresponding exponents.

OpenStudy (anonymous):

Thank You

jimthompson5910 (jim_thompson5910):

You're welcome